OpenStudy (anonymous):

Find the slope-intercept of a line with the following points (-1,2) and (1,-1)

OpenStudy (anonymous):

first find the slope from these points can you???

OpenStudy (anonymous):

M = (-1 - 2) / (1 + 1) M = -3 / 2 M = -1.5?

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

now use point slope form to get the equation

OpenStudy (anonymous):

how do i find x and b?

OpenStudy (anonymous):

using this formula you will get it y-y1 = M (x-x1)

OpenStudy (anonymous):

so y - 2 = -1.5(x + 1) y = -1.5x + 2?

OpenStudy (anonymous):

right , but instead of 2 it will be 0.5

OpenStudy (anonymous):

y = -1.5x + 0.5

OpenStudy (anonymous):

yes this man right

OpenStudy (anonymous):

@rizwan_uet , I'm a girl :P

OpenStudy (anonymous):

yay! why is it .5 instead of 2 though?

OpenStudy (anonymous):

o o i m really sorry

OpenStudy (anonymous):

@Lyubasha

OpenStudy (anonymous):

2 - 1.5 = 0.5 thats why

OpenStudy (anonymous):

@rizwan_uet it's ok , happens all the time :P

OpenStudy (anonymous):

oh i see now